# On Bayes factor functions

### v2

@@ 5.2 Pearson’s $\chi^{2}$ statistics
In this scenario, when $q_{0.5,5}=4.35$, the Bayes factor yields decisive evidence ($BF_{10}(x)<0.01$) against alternative hypotheses for inverse gamma densities centered on standardized effect vectors for which $\sqrt{{\bm{\omega}}^{\prime}{\bm{\omega}}/k}>0.19$. Conversely, when $q_{0.99,5}=15.1$, strong evidence is obtained in favor of the alternative for values of $\sqrt{{\bm{\omega}}^{\prime}{\bm{\omega}}/k}$ in the interval $(0.10,0.14)$.
@@ 6 Analysis of replicated experiments and meta-analyses
To account for both within- and between-study variability, researchers have increasingly relied on hierarchical Bayesian models. These models introduce latent parameters that represent study-level effects and place priors over their distribution. This framework supports coherent posterior inference that reflects uncertainty in nuisance parameters and provides interpretable estimates for quantities of interest such as overall effects, variance components, and predictive distributions for future studies (e.g., Higgins and Spiegelhalter, 2002; Higgins, Thompson and Spiegelhalter, 2009; Gelman et al., 2013).
@@
Recent work by Nikolakopoulos and Ntzoufras (2021) has made this limitation explicit, demonstrating that naively multiplying Bayes factors across studies can lead to misleading inferences. If priors are set independently for each study without regard to a common hierarchical structure or shared effect size, then the resulting Bayes factors may reflect disparate prior beliefs and cannot be meaningfully aggregated. On the other hand, if a common effect is assumed across studies, the marginal distribution of the data from all studies must be obtained by integrating the product of likelihood functions with respect to the prior density on the common effect, rather than integrating over the prior for each likelihood independently. That is,
$\prod_{i=1}\left[\int f(x_{i}{\,|\,}\lambda_{i})\pi(\lambda_{i}{\,|\,}\nu,\tau,n_{i})d\lambda_{i}\right]\neq\int\prod_{i}\left[f(x_{i}{\,|\,}\lambda)\right]\,\pi(\lambda{\,|\,}\nu,\tau)\,d\lambda.$ (15)
In this section, we show how Bayes Factor Functions (BFFs) naturally incorporate features of Bayesian hierarchical modeling by centering prior distributions on standardized effects and explicitly accounting for variability in noncentrality parameters and effect sizes across studies. This framework yields Bayes factors that directly support specific hypotheses about standardized effects. While default values for prior dispersion parameters may be employed, we instead propose method-of-moments estimators for these parameters. When BFFs are constructed using inverse-moment or inverse-gamma priors, the resulting estimator of the dispersion parameter $\nu$ is independent of the standardized effect, allowing a single estimate of $\nu$ to be used to define the entire BFF.
@@ 6.1 Empirical Bayes estimates of $\nu$
$E(\lambda_{i})=\frac{\tau}{\nu/2-1},\quad Var(\lambda_{i})=\frac{\tau^{2}}{(\nu/2-1)^{2}(\nu/2-2)}=\frac{E(\lambda_{i})^{2}}{\nu/2-2}$ (16)
for $\nu>4$. Given a sample of $\{\lambda_{i}\}$, a simple method-of-moments estimate of $\nu$ is thus
@@
In analyzing replicated experiments, the non-centrality parameters of the test statistics are not directly observed. Furthermore, they typically vary across individual experiments $i$ as a function of the sample size, $n_{i}$. Thus, it is necessary to standardize the test statistics before calculating method-of-moments estimates. For example, if $z$ and $t$ tests are based on normal i.i.d. samples, $\lambda_{i}=\sqrt{n_{i}}\omega$, were $\omega$ is the common standardized effect. In that case, dividing the test statistics by $\sqrt{n_{i}}$ produces statistics that share a common non-centrality parameter. Ignoring terms of order $O(1/n_{i})$, we obtain a closed-form estimator for $\nu$ given by:
@@
and $x_{i}$, $i=1,\dots,M$, denote standardized test statistics. For one-sample $z$ and $t$ tests listed in Table 1, $x_{i}=z^{2}_{i}/n_{i}$ or $x_{i}=t_{i}^{2}/n_{i}$, while for the $\chi^{2}$ and $F$ statistics, $x_{i}$ is simply the statistic divided by $n_{i}$. A more formal justification for these estimators is provided in the Supplementary Material.
@@ 6.2 An example: persistence and conscientiousness data
An experiment replicated in the Many Labs 3 project examined the correlation between conscientiousness and persistence. Conscientiousness was measured using two items on the Ten Item Personality Inventory (TIPI; Gosling, Rentfrow and Swann (2003)) and persistence by the time participants spent solving anagrams, some of which were solvable and others not. The study performed in the Many Labs 3 project represented a modification of a study performed initially in De Fruyt, Van de Wiele and Van Heeringen (2000). Table 3 provides the sample correlations and sample sizes collected in 20 replicated studies.
@@
The null hypothesis in this study is that there is no correlation between persistence and conscientiousness measures. To model the sample correlation coefficients, we rescaled Fisher’s z-transformation of the sample correlation coefficient Fisher (1915) and defined
$z_{i}=\frac{\sqrt{n_{i}-3}}{2}\log\left(\frac{1+\hat{\rho}_{i}}{1-\hat{\rho}_{i}}\right).$ (19)
@@
$\qquad z_{i}{\sim}N\left(\lambda_{i},1\right),$ (20)
@@
$\lambda_{i}=\frac{\sqrt{n_{i}-3}}{2}\log\left(\frac{1+\rho_{i}}{1-\rho_{i}}\right)$
@@
Under the null hypothesis, we assume $H_{0}:\mathbf{\rho}=\mathbf{0}$. Under the alternative hypothesis, we define the standardized effect to be
$\omega=\frac{1}{2}\log\left(\frac{1+\rho}{1-\rho}\right)$ (21)
@@
$\lambda_{i}\mid\omega,\nu\sim i(\lambda\mid\tau_{\omega,\nu_{i}},\nu),\qquad\mbox{where}\qquad\tau_{\omega,\nu,i}=\frac{(n_{i}-3)(\nu+1)\omega}{2}.$ (22)
@@
For these data, the method-of-moments estimator of $\nu$ was obtained from (18) using the sample mean and variance of the $z_{i}^{2}/(n_{i}-3)$ values, and equaled $\hat{\nu}_{\text{MOM}}=4.75$
@@
$BF_{10}(\mathbf{z}\mid\omega,\nu)$ $=$ $\frac{\int\cdots\int\prod_{i=1}^{20}n(z_{i}{\,|\,}\lambda_{i},1)i(\lambda_{i}{\,|\,}\tau_{\omega,\nu,i},\nu)d\lambda_{i}}{\prod_{i=1}^{20}n(z_{i}{\,|\,}0,1)}$ (23)
$=$ $\prod_{i=1}^{20}BF_{10}(z_{i}|\tau_{\omega,\nu,i},\nu).$ (24)
@@
As the analysis of EPP in Section 3 suggests, the BFF for $\nu=1$ accumulates evidence more quickly in favor of the null hypothesis in this case than other values of $|nU$. This occurs because the true standardized effect is very likely to less than 0.1, although strong evidence for this conclusion is not provided by individual replications of the experiment.
@@ 7 Discussion
In this article, the performance of several classes of alternative prior densities used to construct BFFs was compared. In the case of true null hypotheses, the asymptotic convergence rates of Bayes factors and BFFs for $z$ and $t$ tests using inverse-moment priors were faster than the convergence rates obtained for the same tests with normal moment, JZS or $g$ priors. Similarly, BFFs based on $\chi^{2}$ and $F$ statistics with inverse gamma priors had better asymptotic convergence rates than those based on gamma, squared Cauchy, and exponential priors on the noncentrality parameters. Expected posterior probability plots generally confirmed the relative performance of these Bayes factors in finite sample settings, except that the JZS and squared Cauchy priors performed similarly (and the squared Cauchy better) for detecting small effect sizes with priors concentrated near the null value. However, the EPP in regions where JZS and squared Cauchy priors perform best occurs when the expected posterior probabilities are less than 0.7, and experimental evidence from a single study is likely to be inconclusive. Inverse moment priors with $\nu=1$ performed comparably to Cauchy distributions for small effect sizes and provided better performance for priors that were less concentrated around the null.
@@
BFFs constructed with normal moment priors ($z$ and $t$ tests) and gamma priors ($\chi^{2}$ and $F$ tests) with moderate values of $r$ performed similarly to inverse-moment and inverse gamma priors for comparable values of $\nu$. Normal moment and gamma priors can also be expressed in closed-form (Datta et al., 2025), making these prior attractive for constructing BFFs to detect moderate to large effect sizes.
BFFs provide a particularly practical framework for synthesizing evidence in meta-analyses and replicated study designs, especially in scenarios where only classical test statistics or $p$-values are available, and full likelihood or design information is lacking. By centering the prior distributions for individual test statistics on noncentrality parameters scaled according to sample size, Bayes factors from separate studies can be coherently aggregated. The resulting cumulative Bayes factor can then be used to construct a BFF that reflects evidence in favor of a range of standardized effects.

### v3

@@ 5.2 Pearson’s $\chi^{2}$ statistics
In this scenario, when $q_{0.5,5}=4.35$, the Bayes factor yields decisive evidence
($BF_{10}(x)<0.01$) against alternative hypotheses for inverse gamma densities centered on standardized effect vectors for which $\sqrt{{\bm{\omega}}^{\prime}{\bm{\omega}}/k}>0.19$. Conversely, when $q_{0.99,5}=15.1$, strong evidence is obtained in favor of the alternative for values of $\sqrt{{\bm{\omega}}^{\prime}{\bm{\omega}}/k}$ in the interval $(0.10,0.14)$.
@@ 6 Analysis of replicated experiments and meta-analyses
To account for both within- and between-study variability, researchers have increasingly relied on hierarchical Bayesian models. These models introduce latent parameters that represent study-level effects and place priors over their distribution. This framework supports coherent posterior inference that reflects uncertainty in nuisance parameters and provides interpretable estimates for quantities of interest such as overall effects, variance components, and predictive distributions for future studies (e.g., Higgins and Spiegelhalter,, 2002; Higgins et al.,, 2009; Gelman et al.,, 2013).
@@
Recent work by Nikolakopoulos and Ntzoufras, (2021) has made this limitation explicit, demonstrating that naively multiplying Bayes factors across studies can lead to misleading inferences. If priors are set independently for each study without regard to a common hierarchical structure or shared effect size, then the resulting Bayes factors may reflect disparate prior beliefs and cannot be meaningfully aggregated. On the other hand, if a common effect is assumed across studies, the marginal distribution of the data from all studies must be obtained by integrating the product of likelihood functions with respect to the prior density on the common effect, rather than integrating over the prior for each likelihood independently. That is,
$\prod_{i=1}\left[\int f(x_{i}{\,|\,}\lambda_{i})\pi(\lambda_{i}{\,|\,}\nu,\tau,n_{i})d\lambda_{i}\right]\neq\int\prod_{i}\left[f(x_{i}{\,|\,}\lambda)\right]\,\pi(\lambda{\,|\,}\nu,\tau)\,d\lambda.$
In this section, we show how BFFs defined with non-local priors naturally incorporate features of Bayesian hierarchical modeling by centering prior distributions on standardized effects and explicitly accounting for variability in noncentrality parameters and effect sizes across studies. This framework yields Bayes factors that directly support specific hypotheses about standardized effects. While default values for prior dispersion parameters may be employed, we instead propose method-of-moments estimators for these parameters. When BFFs are constructed using inverse-moment or inverse-gamma priors, the resulting estimator of the dispersion parameter $\nu$ is independent of the standardized effect, allowing a single estimate of $\nu$ to be used to define the entire BFF.
@@ 6.1 Empirical Bayes estimates of $\nu$
$E(\lambda_{i})=\frac{\tau}{\nu/2-1},\quad Var(\lambda_{i})=\frac{\tau^{2}}{(\nu/2-1)^{2}(\nu/2-2)}=\frac{E(\lambda_{i})^{2}}{\nu/2-2}$
for $\nu>4$. The assumption that $\nu>4$ implies that the variance of the non-centrality parameters is finite. Given a sample of $\{\lambda_{i}\}$, a simple method-of-moments estimate of $\nu$ is thus
@@
In analyzing replicated experiments, the non-centrality parameters of the test statistics are not directly observed. Furthermore, they typically vary across individual experiments $i$ as a function of the sample size, $n_{i}$. Thus, it is necessary to standardize the test statistics before calculating method-of-moments estimates. For example, if $z$ and $t$ tests are based on normal i.i.d. samples, $\lambda_{i}=\sqrt{n_{i}}\omega$, where $\omega$ is the common standardized effect. In that case, dividing the test statistics by $\sqrt{n_{i}}$ produces statistics that share a common non-centrality parameter. Ignoring terms of order $O(1/n_{i})$, we obtain a closed-form estimator for $\nu$ given by:
@@
and $x_{i}$, $i=1,\dots,M$, denote standardized test statistics. For one-sample $z$ and $t$ tests listed in Table 1, $x_{i}=z^{2}_{i}/n_{i}$ or $x_{i}=t_{i}^{2}/n_{i}$, while for the $\chi^{2}$ and $F$ statistics, $x_{i}$ is simply the statistic divided by $n_{i}$. Because $\hat{\nu}_{MOM}>4$, the asymptotic convergence rates cited in Section 3 extend to the resulting aggregated BFFs. A more formal justification for these estimators is provided in the Supplementary Material.
@@ 6.2 An example: persistence and conscientiousness data
An experiment replicated in the Many Labs 3 project (Ebersole et al., (2016)) examined the correlation between conscientiousness and persistence. Conscientiousness was measured using two items on the Ten Item Personality Inventory (TIPI; Gosling et al., (2003)) and persistence by the time participants spent solving anagrams, some of which were solvable and others not. The study performed in the Many Labs 3 project represented a modification of a study performed initially in De Fruyt et al., (2000). Table 2 provides the sample correlations and sample sizes collected in 20 replicated studies.
@@
The null hypothesis in this study is that there is no correlation between persistence and conscientiousness measures. To model the sample correlation coefficients, we rescaled Fisher’s z-transformation of the sample correlation coefficient Fisher, (1915) and defined
$z_{i}=\frac{\sqrt{n_{i}-3}}{2}\log\left(\frac{1+\hat{\rho}_{i}}{1-\hat{\rho}_{i}}\right).$
@@
$\qquad z_{i}{\sim}N\left(\lambda_{i},1\right),$
@@
$\lambda_{i}=\frac{\sqrt{n_{i}-3}}{2}\log\left(\frac{1+\rho_{i}}{1-\rho_{i}}\right).$
@@
Under the null hypothesis, we assume $H_{0}:\ \bm{\rho}=\mathbf{0}$. Under the alternative hypothesis, we define the standardized effect to be
$\omega=\frac{1}{2}\log\left(\frac{1+\rho}{1-\rho}\right)$
@@
$\lambda_{i}\mid\omega,\nu\sim i(\lambda\mid\tau_{\omega,\nu,i},\nu),\qquad\mbox{where}\qquad\tau_{\omega,\nu,i}=\frac{(n_{i}-3)(\nu+1)\omega}{2}.$
@@
For these data, the method-of-moments estimator of $\nu$ was obtained from (5) using the sample mean and variance of the $z_{i}^{2}/(n_{i}-3)$ values and equaled $\hat{\nu}_{\text{MOM}}=4.75$.
@@
$BF_{10}(\mathbf{z}\mid\omega,\nu)$ $=$ $\frac{\int\cdots\int\prod_{i=1}^{20}n(z_{i}{\,|\,}\lambda_{i},1)i(\lambda_{i}{\,|\,}\tau_{\omega,\nu,i},\nu)d\lambda_{i}}{\prod_{i=1}^{20}n(z_{i}{\,|\,}0,1)}$
$=$ $\prod_{i=1}^{20}BF_{10}(z_{i}|\tau_{\omega,\nu,i},\nu).$
@@
As the analysis of $\rm D_{EP}$} in Section 3 suggests, the BFF for $\nu=1$ accumulates evidence more quickly in favor of the null hypothesis in this case than for other values of $\nu$. This occurs because the true standardized effect is very likely to be less than 0.1, although strong evidence for this conclusion is not provided by individual replications of the experiment. Nonetheless, these results suggest that the conservative choice of $\nu=1$ may be preferred for analyzing replicated designs targeted at detecting small standardized effects.
@@ 7 Discussion
In this article, the performance of several classes of alternative prior densities used to construct BFFs was compared. In the case of true null hypotheses, the asymptotic convergence rates of Bayes factors and BFFs for $z$ and $t$ tests using inverse-moment priors were faster than the convergence rates obtained for the same tests with normal moment, JZS or $g$ priors. Similarly, BFFs based on $\chi^{2}$ and $F$ statistics with inverse gamma priors had better asymptotic convergence rates than those based on gamma, squared Cauchy, and exponential priors on the noncentrality parameters. Expected posterior probability plots generally confirmed the relative performance of these Bayes factors in finite sample settings, except that the JZS and squared Cauchy priors performed similarly (and the squared Cauchy better) for detecting small effect sizes with priors concentrated near the null value. However, the $\rm D_{EP}$} in regions where JZS and squared Cauchy priors perform best occurs when the expected posterior probabilities are less than 0.7 and experimental evidence from a single study is likely to be inconclusive. Inverse moment priors with $\nu=1$ performed comparably to Cauchy distributions for small effect sizes and provided better performance for priors that were less concentrated around the null.
@@
BFFs constructed with normal moment priors ($z$ and $t$ tests) and gamma priors ($\chi^{2}$ and $F$ tests) with moderate values of $r$ performed similarly to inverse-moment and inverse gamma priors for comparable values of $\nu$. Normal moment and gamma priors can also be expressed in closed-form (Datta et al.,, 2025), making these prior attractive for constructing BFFs to detect moderate to large effect sizes.
BFFs provide a particularly practical framework for synthesizing evidence in meta-analyses and replicated study designs, especially in scenarios where only classical test statistics or $p$-values are available, or when full likelihood or design information is unavailable. By centering the prior distributions for individual test statistics on noncentrality parameters scaled according to sample size, Bayes factors from separate studies can be coherently aggregated. The resulting cumulative Bayes factor can then be used to construct a BFF that reflects evidence in favor of a range of standardized effects.